Need help with a percent problem please!
If 2.4 million dollars is spent annually in CA on frozen pizza, how much more is spent on pepperoni pizzas than sausage pizzas each year? 31% pepperoni pizzas and 20% for sausage pizza

Found 2 solutions by edjones, davethejackal:
Answer by edjones(8007)

2.4*0.31=744,000
2.4*0.20=480,000
========>264,000 more is spent on pepperoni pizzas than sausage pizzas each year
.
Ed

Answer by davethejackal(28)

100% of the pizza market is $2.4million or $2,400,000.
This si 100% so it is 100 x more than 1% of the total market.
In other words 1% of the market is 1/100 of the total market ...
or $2,400,000/100 = $24,000
The pepperoni market is 31%, which is 31 x 1% of the market = 31 x 24000 = $744,000
Sausage pizza is 20%, which is 20% 1% of the market = 20 x 24000 = $480,000
So, the difference between the two is 744,000 - 480,000 = $264,000